OpenStudy (anonymous):

The sum of the even integers between 21 and 81

OpenStudy (anonymous):

22,24,26........80 ita an arithmetic sequence :)

OpenStudy (anonymous):

how do you get the sum?

OpenStudy (godlovesme):

Srrry idk =0

hartnn (hartnn):

use the formula : \(\Large S_n = n/2 (2a+(n-1)d)\) familiar with this formula?

hartnn (hartnn):

n is the number of terms that you need to find out first a = 1st term d = common difference

OpenStudy (anonymous):

thanks !

OpenStudy (perl):

you can use the sequence a(n) = 2n + 20

OpenStudy (perl):

a(1) = 2*1 + 20 = 22 a(30) = 2*30 + 20 = 80